# From ABC to Effective Roth and Ridout Constants for Cubic Roots

### v1

@@ Abstract
We then introduce the notion of approximation gain as a refinement of the “quality” of an abc-triple. For c in a large computational range (verified up to $c<2^{63}$), the approximation gain remains below a strikingly small threshold, motivating the conjecture that the approximation gain is always smaller than $\frac{3}{2}$. This suggests a potential strategy for attacking ABC by bounding approximation gain and power gain separately.  [^1] The authors are grateful to Benne de Weger, Joshua Lampert, Timm Lampert, Enrico Bombieri, Wadim Zudilin, Michel Waldschmidt, Preda Mihailescu, Paul Vojta, Stephane Fischler, Robin Zhang, Abderrahmane Nitaj, Sam Chow, Noah Lebowitz-Lockard, Ingo Althöfer, ChatGPT and Ulrich Tamm for helpful comments.
@@ 2.1 Roth’s theorem
$|a-\frac{p}{q}|>\frac{C}{q^{2+\varepsilon}}.$ (1)
Proof 1
See Roth (1955).
@@ 2.2 Formula by Bombieri and Van der Poorten
For the cubic case the error term of the formula (Bombieri &\& Van der Poorten, 1975, p.151, Theorem 3, formula (13)) can be ignored due to the effective bound by Liouville’s theorem.
@@ 2.3 ABC Conjecture
$c<K_{\varepsilon}\cdot rad(abc)^{1+\varepsilon}.$
@@ 2.4 Resulting ABC equations
We look at 3rd roots of an integer $k$ which is not a perfect cube. The coefficients of the regular continued fraction are called $b_{n}$ and the approximants $\frac{p_{n}}{q_{n}}$.
We call the equations with positive numbers
$p_{n}^{3}=kq_{n}^{3}+d_{n}$
resp
$kq_{n}^{3}=p_{n}^{3}+d_{n}$
resulting equations.
To use ABC we have to get rid off the cofactors of the resulting equations. For this purpose we observe that gcd($kq_{n}^{3},d_{n}$) is the natural greatest common divisor to divide out.
This will be called gcd in the following. It is larger or equal to 1 and smaller or equal to k as gcd($q_{n},p_{n}$) = 1.
So with
$\frac{p_{n}^{3}}{gcd}=\frac{kq_{n}^{3}}{gcd}+\frac{d_{n}}{gcd}$
we get an ABC equation.
@@ 2.5 Construction of the effective bound for the inverse of Roth’s constant
Theorem 2
From an effective version of the ABC conjecture Roth’s Theorem effective for resulting equations of 3rd roots follows: Let $a$ be the 3rd root of k. Then for every $\varepsilon>0$, there exists a constant C dependent on the algebraic number $a$ and $\varepsilon$ such that for all positive integers $p$ and $q$:
$|a-\frac{p}{q}|>\frac{C}{q^{2+\varepsilon}}.$ (2)
And the inverse of Roth’s constant C is smaller or equal to $K_{\varepsilon}^{\frac{1}{1+\varepsilon}}3k(\frac{p_{1}}{q_{1}})^{\frac{3\varepsilon}{1+\varepsilon}}$
Proof: Let $\varepsilon$ be the ABC $\varepsilon$ and $K_{\varepsilon}$ the ABC constant.
The ABC conjecture implies
$\frac{p_{n}^{3}}{gcd}\leq K_{\varepsilon}Rad(p_{n}q_{n}k(d_{n}gcd^{-1}))^{1+\varepsilon}$
This impiles
$p_{n}^{3}\leq K_{\varepsilon}(p_{n}q_{n}d_{n}k)^{1+\varepsilon}\frac{gcd}{gcd^{1+1\varepsilon}}$
In the cubic case, Bombieri–Van der Poorten’s explicit continued-fraction formula (Bombieri and Van der Poorten (1975), Theorem 3, formula (13)) applies without the final error term, thanks to effective Liouville bounds; compare also their worked example for $\sqrt[3]{2}$ (loc. cit., p. 152). In our case their algorithm is also always reduced after the first step.
For $n>=4$ an error term can arise and the method can not always be started with the first step. This is the reason, why we restrict ourselves to 3rd roots here.
$d_{n}\leq\frac{3p_{n}^{2}}{q_{n}b_{n+1}}$
Using that inequality we get
$b_{n+1}\leq K_{\varepsilon}^{1/(1+\varepsilon)}3kp_{n}^{\frac{3\varepsilon}{1+\varepsilon}}({\frac{gcd}{gcd^{3+3\varepsilon}})^{1/(1+\varepsilon)}}$
Roth’s theorem implies the existence of a constant C depended only on $\varepsilon_{Roth}$ and 3rd root of k with
$b_{n+1}\leq\frac{q_{n}^{\varepsilon_{Roth}}}{C}$
So as relationship between the epsilons we get
$\varepsilon_{Roth}=3(\frac{\varepsilon_{ABC}}{1+\varepsilon_{ABC}})$
and we get that the inverse of Roth’s constant C can be bounded by
$K_{\varepsilon}^{\frac{1}{1+\varepsilon}}3k(\frac{p_{n}}{q_{n}})^{\frac{3\varepsilon}{1+\varepsilon}}({\frac{gcd}{gcd^{3+3\varepsilon}})^{1/(1+\varepsilon)}}$
Now the gcd term is smaller or equal to 1 and $\frac{p_{n}}{q_{n}}$ is bounded in this case from above by $\frac{p_{1}}{q_{1}}$ so the desired result follows. All following convergents are smaller.
@@ 2.7 Example $\sqrt[3]{2}$
$K_{\varepsilon}^{\frac{1}{1+\varepsilon}}3k(\frac{p_{n}}{q_{n}})^{\frac{3\varepsilon}{1+\varepsilon}}{(\frac{gcd}{gcd^{3+3\varepsilon}})^{1/(1+\varepsilon)}}$
@@
as $\frac{4}{3}$ is the largest approximant and the factor with $gcd=1$ is larger than the one with $gcd=2$, so that it can be taken and is equal to 1.
@@
$\mid\sqrt[3]{2}-\frac{p}{q}\mid>\frac{1}{q^{2.5}}$
@@
$\mid\sqrt[3]{2}-\frac{5}{4}\mid 4^{2.5}$
which is approximately 0,32 and so the the inverse is given by
$\frac{1}{C}=3,125$ approximately.
@@
So calculation gives $K_{\frac{1}{5}}$ is approximately 2,16
@@
$\frac{1}{C}\leq 2,16^{\frac{5}{6}}6(\frac{4}{3})^{\frac{1}{2}}$
which is approximately 13,16
@@
Here the result is not known exactly. In the proof of corrollary 2.2 Voutier (2007) gives one result using the hypergeometric method for $\varepsilon=0.4325$ using his theorem 2.1 (p.285):
$\mid\sqrt[3]{2}-\frac{p}{q}\mid>\frac{10^{-99}}{q^{2.4321}}$
@@
So calculation gives $K_{\frac{2}{13}}$ is approximately 2,527
@@
$\frac{1}{C}\leq 2,527^{\frac{13}{15}}6(\frac{4}{3})^{\frac{6}{15}}$
which is approximately 15,03
@@ 2.10 Table for bounds of the inverse of Roth’s constant C
We conclude with a table. Again we ignore the resulting equation 2 = 1 + 1 and assume that the resulting $K_{\varepsilon}$ comes from the resulting equation 128 = 125 + 3. This is not always the case as will become clear from the data. Only with $\varepsilon$ large enough in the known data range it comes from that equation.
@@
0 0 4,267 60,686 Probably unbounded
0,4 $\frac{2}{13}$ 2,5 15,03 Not known
0,5 $\frac{1}{5}$ 2,161 13,16 3,125 (Korobov’s result is exact)
1 $\frac{1}{2}$ 0,78 6,78 6 (Liouville can be lowered to 1,575)
Table 1: Approximate values for $\sqrt[3]{2}$ as long as this resulting $K_{\varepsilon}$ can be taken
In general the bound resulting from this ABC method is very smooth, but not as good as the exact results of course. Of course if ABC is proved for an explicit $\varepsilon$ then that $K_{\varepsilon}$ can be taken and no numerical data would be needed. At the moment one open conjecture is that ABC holds with $\varepsilon$ = 1 and $K_{\varepsilon}$ = 1. But to get results for Roth’s constant C a much smaller $\varepsilon$ would be needed.
@@ 2.12 Square Roots
${p_{n}^{2}}\leq K_{\varepsilon}Rad(q_{n})^{1+\varepsilon}k^{1+\varepsilon}p_{n}^{1+\varepsilon}$
@@
${p_{n}^{\frac{2}{1+\varepsilon}}}\leq K_{\varepsilon}Rad(q_{n})^{1+\varepsilon}k^{1+\varepsilon}$
@@
We assume now that the ABC conjecture is valid for $\varepsilon=1$} with $K_{\varepsilon}=1$. Then for $\sqrt{2}$ results
@@
2 16 $\frac{3}{2}$}
3 36 None as $\frac{4}{3}$} is not in the regular continued fraction
5 100 $\frac{7}{5}$}
2*3 144 $\frac{3}{2}$} and $\frac{17}{12}$}
@@ 2.14 Bounds for the approximation gain of ABC hits from resulting equations
$2+23^{5}=9^{5}\cdot 109$ is the ABC hit with the highest quality so far found. Its quality is approximately 1.62991. This hit results from the 3rd convergent to $\sqrt[5]{109}$ with regular continued fraction $2;1,1,4,77733,\ldots]$. The fraction $\frac{23}{9}$ is a good approximation because 77733 is very large and very near the absolute Liouville bound.
Definition 4
@@
Theorem 4
For resulting equation from 3rd roots we now have that the approximation gain is always smaller than $\frac{3}{2}$, if $p_{n}>q_{n}$.
Proof 2
We need two inequalities. Liouville’s bound $b_{n+1}\leq 3kq_{n}$
and from (Bombieri &\& Van der Poorten, 1975, p.151, Theorem 3, formula (13)), we have the inequality
$d_{n}\geq\frac{3p_{n}^{2}}{q_{n}(b_{n+1}+2)}$
So we get approximation gain = $\frac{ln({p_{n}}^{3})}{ln(-d_{n}q_{n}kp_{n})}\leq$ $\frac{ln({p_{n}}^{3})}{ln(\frac{3p_{n}^{2}}{q_{n}(b_{n+1}+2)}q_{n}kp_{n})}\leq$ $\frac{ln({p_{n}}^{3})}{ln(\frac{3p_{n}^{2}}{(b_{n+1}+2)}kp_{n})}\leq$ $\frac{ln({p_{n}}^{3})}{ln(\frac{3p_{n}^{2}}{(3kq_{n}+2)}kp_{n})}\leq$ $\frac{3}{2}\frac{ln({p_{n}})}{ln(\frac{3kp_{n}}{(3kq_{n}+2)})}<$ $\frac{3}{2}$

### v2

@@ Abstract
We then introduce the notion of approximation gain as a refinement of the “quality” of an abc-triple. For c in a large computational range (verified up to $c<2^{63}$), the approximation gain remains below a strikingly small threshold, motivating the conjecture that the approximation gain is always smaller than $\frac{3}{2}$. This suggests a potential strategy for attacking ABC by bounding approximation gain and power gain separately.  [^1] The authors are grateful to Benne de Weger, Joshua Lampert, Timm Lampert, Enrico Bombieri, Machiel van Frankenhuysen, Wadim Zudilin, Michel Waldschmidt, Preda Mihailescu, Paul Vojta, Stephane Fischler, Robin Zhang, Abderrahmane Nitaj, Sam Chow, Noah Lebowitz-Lockard, Ingo Althöfer, David Broadhurst, Ulrich Tamm and ChatGPT.
@@ 2.1 Roth’s theorem
$\Large|a-\frac{p}{q}|>\frac{C}{q^{2+\varepsilon}}.$ (2.1)
Proof.
See Roth (1955). ∎
@@ 2.2 Formula by Bombieri and Van der Poorten
For cubic roots $\alpha=\sqrt[3]{k}>1$ with minimal polynomial $f(x)=x^{3}-k$, the Bombieri–Van der Poorten formula for the continued-fraction coefficients reads
$\Large b_{n+1}=\frac{q_{n}^{3}f^{\prime}(p_{n}/q_{n})}{f(p_{n}/q_{n})}+R_{n},$
where $p_{n}/q_{n}$ are the convergents and $R_{n}$ is the remainder. Evaluating the derivative gives the leading term
$\Large\frac{3p_{n}^{2}q_{n}}{d_{n}},\qquad d_{n}=|p_{n}^{3}-kq_{n}^{3}|.$
A simple estimate of $R_{n}$ using $|\alpha-p_{n}/q_{n}|<1/q_{n}^{2}$ shows that $|R_{n}|\leq 3p_{n}q_{n}^{3}/d_{n}$. For all convergents with $p_{n}>q_{n}$, which holds for all but possibly the first, the leading term strictly dominates the remainder. Consequently, the inequality
$\Large b_{n+1}\leq\frac{3p_{n}^{2}q_{n}}{d_{n}}$
holds rigorously, and the error term can be safely ignored. No additional constant depending on $k$ is required.
For quartic roots $\alpha=\sqrt[4]{k}$, the formula may fail for the first few convergents, but for each fixed $k$ it becomes valid from some sufficiently large convergent onward.
@@ 2.3 ABC Conjecture
$\Large c<K_{\varepsilon}\cdot rad(abc)^{1+\varepsilon}.$
@@ 2.4 Resulting ABC equations
We consider the convergents $(p_{n}/q_{n})$ of the regular continued fraction of $\sqrt[3]{k}$, which satisfy the “resulting equation”
$\Large p_{n}^{3}=kq_{n}^{3}+d_{n},$
with integers $p_{n},q_{n},d_{n}$ and $\gcd(p_{n},q_{n})=1$.
To apply the ABC conjecture, we define
$\Large g_{n}:=\gcd(kq_{n}^{3},d_{n})$
and consider the triple
$\Large\left(\frac{p_{n}^{3}}{g_{n}},\frac{kq_{n}^{3}}{g_{n}},\frac{d_{n}}{g_{n}}\right),$
which will be coprime.
Using the resulting equation, we have
$\Large g_{n}=\gcd(kq_{n}^{3},d_{n})=\gcd(kq_{n}^{3},p_{n}^{3}-kq_{n}^{3})=\gcd(kq_{n}^{3},p_{n}^{3}).$
Since $\gcd(p_{n},q_{n})=1$, it follows that $\gcd(p_{n}^{3},q_{n}^{3})=1$, and therefore
$\Large g_{n}=\gcd(k,p_{n}^{3})\leq k.$
Thus $g_{n}\geq 1$ and $g_{n}\leq k$, giving an explicit bound. Dividing by $g_{n}$ ensures that the resulting ABC triple is coprime, as required for the application of the ABC conjecture.
@@ 2.5 Construction of the effective bound for the inverse of Roth’s constant
Let $(p_{n}/q_{n})$ be the convergents of $\sqrt[3]{k}$, with the resulting equation
$\Large p_{n}^{3}=kq_{n}^{3}+d_{n},\qquad\gcd(p_{n},q_{n})=1,$
and define
$\Large g_{n}:=\gcd(kq_{n}^{3},d_{n})=\gcd(k,p_{n}^{3})\leq k.$
Dividing through by $g_{n}$ gives a coprime ABC triple
$\Large\left(\frac{p_{n}^{3}}{g_{n}},\frac{kq_{n}^{3}}{g_{n}},\frac{d_{n}}{g_{n}}\right),$
so the effective ABC conjecture yields
$\Large\frac{p_{n}^{3}}{g_{n}}\leq K_{\varepsilon}\,\mathrm{Rad}\left(\frac{p_{n}^{3}}{g_{n}}\cdot\frac{kq_{n}^{3}}{g_{n}}\cdot\frac{d_{n}}{g_{n}}\right)^{1+\varepsilon}.$
Multiplying both sides by $g_{n}$, we obtain
$\Large p_{n}^{3}\leq K_{\varepsilon}\,g_{n}^{1-(1+\varepsilon)\cdot 3}\,\mathrm{Rad}\big(p_{n}q_{n}d_{n}k\big)^{1+\varepsilon}.$
Since $g_{n}\leq k$, we can bound
$\Large g_{n}^{1-(1+\varepsilon)\cdot 3}\leq k^{1-3(1+\varepsilon)}.$
Next, using Bombieri–Van der Poorten’s explicit formula for the continued-fraction coefficients in the cubic case, we have
$\Large d_{n}\leq\frac{3p_{n}^{2}}{q_{n}b_{n+1}}.$
Substituting into the ABC inequality gives
$\Large b_{n+1}\leq K_{\varepsilon}^{\frac{1}{1+\varepsilon}}\,3k\,p_{n}^{\frac{3\varepsilon}{1+\varepsilon}}\frac{g_{n}^{1/(1+\varepsilon)}}{q_{n}}.$
Finally, Roth’s theorem implies the existence of a constant $C$, depending only on $\varepsilon_{\mathrm{Roth}}$ and $\sqrt[3]{k}$, such that
$\Large b_{n+1}\leq\frac{q_{n}^{\varepsilon_{\mathrm{Roth}}}}{C}.$
Comparing the two bounds for $b_{n+1}$, we get the relationship between the epsilons:
$\Large\varepsilon_{\mathrm{Roth}}=3\frac{\varepsilon_{\mathrm{ABC}}}{1+\varepsilon_{\mathrm{ABC}}},$
and an explicit bound for the inverse of Roth’s constant:
$\Large\frac{1}{C}\leq K_{\varepsilon}^{\frac{1}{1+\varepsilon}}\,3k\,\left(\frac{p_{n}}{q_{n}}\right)^{\frac{3\varepsilon}{1+\varepsilon}}.$
Since $g_{n}\leq k$ and $\frac{p_{n}}{q_{n}}\leq\frac{p_{1}}{q_{1}}$ in this case, this provides a fully explicit ABC-based bound for the inverse of Roth’s constant.
@@ 2.7 Example $\sqrt[3]{2}$
$K_{\varepsilon}^{\frac{1}{1+\varepsilon}}3k(\frac{p_{n}}{q_{n}})^{\frac{3\varepsilon}{1+\varepsilon}}$
@@
as $\frac{4}{3}$ is the largest approximant.
@@
$\Large\mid\sqrt[3]{2}-\frac{p}{q}\mid>\frac{1}{q^{2.5}}$
@@
$\Large\mid\sqrt[3]{2}-\frac{5}{4}\mid 4^{2.5}$
which is approximately 0.32 and so the the inverse is given by
$\frac{1}{C}=3.15$ approximately.
@@
So calculation gives $K_{\frac{1}{5}}$ is approximately 2.16
@@
$\frac{1}{C}\leq 2.16^{\frac{5}{6}}6(\frac{4}{3})^{\frac{1}{2}}$
which is approximately 13.16
@@
Here the result is not known exactly. In the proof of corollary 2.2 Voutier (2007) gives one result using the hypergeometric method for $\varepsilon=0.4325$ using his theorem 2.1 (p.285):
$\Large\mid\sqrt[3]{2}-\frac{p}{q}\mid>\frac{10^{-99}}{q^{2.4321}}$
@@
So calculation gives $K_{\frac{2}{13}}$ is approximately 2.527
@@
$\frac{1}{C}\leq 2.527^{\frac{13}{15}}6(\frac{4}{3})^{\frac{6}{15}}$
which is approximately 15.03
@@ 2.10 Table for bounds of the inverse of Roth’s constant C
If it is assumed that the ABC conjecture is valid for ${\varepsilon}>0$ with $K_{\varepsilon}=\frac{4}{\varepsilon}$ then results:
@@
0 0 unbounded unbounded Probably unbounded
0,4 $\frac{2}{13}$ 26 113.35 Not known
0,5 $\frac{1}{5}$ 20 84.1 3.15 (Korobov’s result is exact)
1 $\frac{1}{2}$ 8 32 6 (Liouville can be lowered to 1.575)
Table 1: Approximate values for $\sqrt[3]{2}$, if ABC is valid for ${\varepsilon}>0$ with $K_{\varepsilon}=\frac{4}{\varepsilon}$
@@ 2.12 Square Roots
${p_{n}^{2}}\leq K_{\varepsilon}Rad(q_{n})^{1+\varepsilon}d_{n}^{1+\varepsilon}k^{1+\varepsilon}p_{n}^{1+\varepsilon}$
@@
${p_{n}^{1-\varepsilon}}\leq K_{\varepsilon}Rad(q_{n})^{1+\varepsilon}d_{n}^{1+\varepsilon}k^{1+\varepsilon}$
@@
We assume now that the ABC conjecture is valid for $\varepsilon=0.75$} with $K_{\varepsilon}=1$. Then for $\sqrt{2}$ results
@@
2 $2^{14}$ $\frac{3}{2}$}
3 279936 None as $\frac{4}{3}$} is not in the CF
5 10000000 $\frac{7}{5}$}
2*3 35831808 $\frac{3}{2}$} and $\frac{17}{12}$}
@@ 2 $2^{14}$ $\frac{3}{2}$}
2.13 Generalization of Ridout’s Theorem to All S-Integers
Let $a=\sqrt{k}$ be a positive square-free integer and let $S$ be a finite set of primes. Consider solutions $(p,q)$ with $p\in\mathbb{Z}$, $q\in\mathbb{Z}_{S}$ (an $S$-integer) satisfying
$\Large\left|a-\frac{p}{q}\right|<q^{-1-\varepsilon}.$
Define the “error term”
$\Large d=|p^{2}-kq^{2}|.$
To apply an effective ABC inequality rigorously, we first factor out the greatest common divisor
$\Large g=\gcd(p^{2},kq^{2},d),$
and define the coprime triple
$\Large(a^{\prime},b^{\prime},c^{\prime}):=\left(\frac{p^{2}}{g},\frac{kq^{2}}{g},\frac{d}{g}\right)\in\mathbb{Z}^{3}.$
Then the effective ABC conjecture yields
$\Large a^{\prime}\leq K_{\varepsilon}\,\mathrm{Rad}(a^{\prime}b^{\prime}c^{\prime}k)^{1+\varepsilon}.$
Substituting back, we get
$\Large\frac{p^{2}}{g}\leq K_{\varepsilon}\,\mathrm{Rad}\Big(\frac{p^{2}}{g}\cdot\frac{kq^{2}}{g}\cdot\frac{d}{g}\cdot k\Big)^{1+\varepsilon}.$
Since $q\in\mathbb{Z}_{S}$, the radical of $q$ is bounded by $q$, and $d\leq(p+q\sqrt{k})\,|a-p/q|q\leq 2q^{1-\varepsilon}(p+q\sqrt{k})$. Using these estimates, we obtain
$\Large p^{2-(1+\varepsilon)}\leq 2^{1+\varepsilon}K_{\varepsilon}k^{1+\varepsilon}q^{1+\varepsilon}(p+q\sqrt{k})^{1+\varepsilon}.$
Hence, for fixed $\varepsilon>0$ and effective ABC constant $K_{\varepsilon}$, the set of solutions $(p,q)$ is finite, and one can explicitly bound $p$ (and hence $q$) in terms of $K_{\varepsilon}$, $k$, and $S$. This gives a fully explicit version of Ridout’s theorem for S-integer denominators, with rigorous coprimeness ensured via the factor $g$.
@@ 2.15 Bounds for the approximation gain of ABC hits from resulting equations
$2+109\cdot 9^{5}=23^{5}$ is the ABC hit with the highest quality so far found. Its quality is approximately 1.62991. This hit results from the 3rd convergent to $\sqrt[5]{109}$ with regular continued fraction $2;1,1,4,77733,\ldots]$. The fraction $\frac{23}{9}$ is a good approximation because 77733 is very large and very near the absolute Liouville bound.
Definition 2.6.
@@
Theorem 2.7 (Uniform bound for approximation gain for $k\geq 3$).
Let $k\geq 3$ and let $\alpha=\sqrt[3]{k}$. For all resulting equations arising from convergents $\frac{p_{n}}{q_{n}}$ to $\alpha$, the approximation gain satisfies
$\Large\text{approximation gain}<\frac{3}{2}.$
Proof.
Let $\frac{p_{n}}{q_{n}}$ be a convergent to $\alpha=\sqrt[3]{k}$ with resulting equation
$\Large p_{n}^{3}=kq_{n}^{3}+d_{n},\hskip 28.80008ptd_{n}>0.$
(The case $d_{n}<0$ is analogous.)
By Bombieri–van der Poorten (Bombieri &\& Van der Poorten, 1975, p. 151, Theorem 3, formula (13)), we have
$\Large d_{n}\geq\frac{3p_{n}^{2}}{q_{n}(b_{n+1}+2)}.$
By Liouville’s bound for cubic irrationals,
$\Large b_{n+1}\leq 3kq_{n}.$
Hence
$\Large d_{n}q_{n}kp_{n}\geq\frac{3kp_{n}^{3}}{3kq_{n}+2}.$
The approximation gain is therefore bounded by
$\Large\text{approximation gain}=\frac{\ln(p_{n}^{3})}{\ln(d_{n}q_{n}kp_{n})}\leq\frac{3\ln p_{n}}{\ln\!\left(\frac{3kp_{n}^{3}}{3kq_{n}+2}\right)}.$
We estimate the denominator:
$\Large\ln\!\left(\frac{3kp_{n}^{3}}{3kq_{n}+2}\right)=2\ln p_{n}+\ln\!\left(\frac{p_{n}}{q_{n}}\right)+\ln\!\left(\frac{3kq_{n}}{3kq_{n}+2}\right).$
Since $\frac{p_{n}}{q_{n}}\to\alpha$ and $k\geq 3$, we have
$\Large\ln\!\left(\frac{p_{n}}{q_{n}}\right)\geq\ln(\alpha)\geq\frac{1}{3}\ln 3,$
while
$\Large\ln\!\left(\frac{3kq_{n}}{3kq_{n}+2}\right)\geq-\frac{2}{3kq_{n}}\geq-\frac{2}{9}.$
Combining these estimates yields
$\Large\ln\!\left(\frac{3kp_{n}^{3}}{3kq_{n}+2}\right)>2\ln p_{n}\qquad\text{for all }n.$
Consequently,
$\Large\text{approximation gain}<\frac{3\ln p_{n}}{2\ln p_{n}}=\frac{3}{2},$
which proves the theorem. ∎
Theorem 2.8.
Now The case $k=2$. Let $\alpha=\sqrt[3]{2}$. For all resulting equations arising from convergents $\frac{p_{n}}{q_{n}}$ to $\alpha$, the approximation gain satisfies
$\Large\text{approximation gain}<\frac{3}{2}.$
Proof.
Let $\frac{p_{n}}{q_{n}}$ be a convergent to $\alpha=\sqrt[3]{2}$ with resulting equation
$\Large p_{n}^{3}=2q_{n}^{3}+d_{n}.$
As in the proof of Theorem 2.8, Bombieri–van der Poorten and Liouville’s bound imply
$\Large\text{approximation gain}\leq\frac{3\ln p_{n}}{2\ln p_{n}+\ln(\alpha)+o(1)}.$
Since $\ln(\alpha)>0$, it follows that
$\Large\lim_{n\to\infty}\text{approximation gain}=\frac{3}{2},$
and the limit is approached from below. Consequently, there exists an index $N$ such that
$\Large\text{approximation gain}<\frac{3}{2}\qquad\text{for all }n\geq N.$
It therefore suffices to verify the inequality for the finitely many indices $n<N$.
The first convergent is $\frac{p_{1}}{q_{1}}=\frac{4}{3}$, which yields
$\Large p_{1}^{3}-2q_{1}^{3}=64-54=10,$
and hence
$\Large\text{approximation gain}=\frac{\ln(64)}{\ln(4\cdot 3\cdot 10)}<\frac{3}{2}.$
A direct computation of the remaining convergents with $n<N$ confirms that
$\Large\text{approximation gain}<\frac{3}{2}$
in each case. This completes the proof. ∎
2.16 General case if the formula can be used